### v1

@@ 4.1 Detection of tidal structures
Based on the matched filtering technique, we constructed S/N maps for all clusters in our sample (see Fig. 3). The maps are presented in a projected coordinate system centered on each cluster, with axes defined as $\Delta\mathrm{RA}\times\cos(\mathrm{Dec})$ and $\Delta\mathrm{Dec}$, ensuring consistent spatial analysis across different celestial positions. The cluster centers were adopted from the Galactic globular cluster database 333people.smp.uq.edu.au/HolgerBaumgardt/globular/, which is extensively used in this study. This database, developed by Baumgardt and collaborators, provides a comprehensive set of physical and dynamical parameters for Galactic globular clusters. These parameters were derived by combining a compilation of observational data, including ground-based radial velocities, Gaia DR3 proper motions, and HST-based stellar mass functions, with literature data, and by performing N-body modeling and orbital integrations (see, e.g., 2017MNRAS.464.2174B; 2018MNRAS.478.1520B; 2021MNRAS.505.5978V; 2021MNRAS.505.5957B; 2023MNRAS.521.3991B, etc.). Further details can be found in the database.
@@
We adopted a uniform spatial resolution of $0.02^{\circ}$ ($1\arcmin.2$) for all clusters except for Whiting 1. This cluster is a relatively small-size system, with a King-model tidal radius of only $0\arcmin.89$ according to the Harris catalog (1996AJ....112.1487H 1996AJ....112.1487H, 2010 edition), and it lies at a heliocentric distance of $30.59$ kpc (2021MNRAS.505.5957B). As a result, its projected angular extent on the sky is limited. To better resolve its structural features, we adopted a finer spatial resolution of $0.01^{\circ}$ ($0\arcmin.6$) for this target. In addition, we applied Gaussian smoothing to the matched-filter S/N maps to enhance the visibility of extended structures while suppressing small-scale noise. Because the clusters in our sample differ substantially in distance, apparent angular size, and the spatial distribution of their member stars, a single smoothing scale is not appropriate. We therefore adopted $\sigma$ = 1 pixel as the default smoothing kernel. The kernel size was then adjusted through systematic visual inspection of the resulting S/N maps to identify the smallest kernel that effectively reduces noise without oversmoothing genuine features. Such inspections help effectively reduce the risk of smoothing out real features or missing potential tidal structures from insufficient smoothing. In practice, the smoothing kernels adopted for our sample range from $\sigma$=1 to 5 pixels, depending on the cluster’s properties.
@@
To investigate the properties of extra-tidal features among our sample clusters, we analyzed the matched-filtered S/N maps and adopted a classification scheme similar to that of 2020A&A...637L...2P. Specifically, we overlaid two reference tidal radii on each map: the King tidal radius (${r_{\rm t}}^{\rm King}$) taken from the Harris catalog (1996AJ....112.1487H 1996AJ....112.1487H, 2010 edition), which is derived from surface density fitting under the King model (1962AJ.....67..471K), and the dynamical tidal radius (${r_{\rm t}}^{\rm dyn}$) from the globular cluster database 444people.smp.uq.edu.au/HolgerBaumgardt/globular/parameter.html, which represents the present-day limiting radius $r_{\rm L}(F)$ calculated following the expression of 2013ApJ...764..124W:
@@
where $R_{\rm GC}$ is the current Galactocentric distance of the cluster, and $R_{\rm Apo}$ and $R_{\rm Peri}$ denote the apocenter (maximum) and pericenter (minimum) distances along its orbit, respectively. The coefficients are $a=0.17\pm 0.03$ and $b=4.1\pm 0.2$ (2013ApJ...764..124W). Compared to the King tidal radius, the dynamical tidal radius provides a physically motivated boundary that accounts for orbital dynamics and gravitational interactions with the Galactic potential. All clusters in our sample are classified into three groups according to their structural properties with respect to $r_{\rm t}^{\rm King}$ and ${r_{\rm t}}^{\rm dyn}$:
  1. G1:
    clusters that exhibit coherent, tail-like extensions aligned along a preferred direction, with detectable overdensities (typically reaching $\mathrm{S/N}\geq 1\sigma$) that extend beyond the dynamical tidal radius ($r_{\rm s}>r_{\rm t}^{\rm dyn}$);
  2. G2:
    clusters that exhibit overdensities with $\mathrm{S/N}\geq 3\sigma$ extending beyond the King tidal radius ($r_{\rm s}>r_{\rm t}^{\rm King}$), while being overall confined within the dynamical tidal radius ($r_{\rm s}<r_{\rm t}^{\rm dyn}$). These extensions typically appear as diffuse envelopes, irregular overdensities, or clumpy halo-like features, and may in some cases extend beyond $r_{\rm t}^{\rm dyn}$. In addition, clusters that display tail-like structures beyond $r_{\rm t}^{\rm King}$ but still remain fully contained within $r_{\rm t}^{\rm dyn}$ are also included in this group;
  3. G3:
    clusters that show no overdensities with $\mathrm{S/N}\geq 3\sigma$ beyond the King tidal radius ($r_{\rm s}<r_{\rm t}^{\rm King}$), and whose outer stellar density profiles are consistent with the truncation expected from a King model.
@@ 4.2 Dependence of tidal features on intrinsic properties
To investigate whether any internal parameters can effectively predict the presence or morphology of tidal structures in globular clusters, we selected a representative set of physical quantities that are commonly associated with the structural and dynamical evolution of clusters. These include the present-day cluster mass ($\log M$), central escape velocity ($v_{\rm esc}$), central concentration ($c$), ratio of half-mass radius to dynamical tidal radius ($r_{\rm h,m}/r_{\rm t}^{\rm dyn}$), mass-loss fraction ($M_{\rm dis}/M_{\rm init}$), metallicity ($\rm[Fe/H]$), and relaxation-age ratio ($\log({\rm age}/{T}_{\rm rh})$). Except for $c$ and $\rm[Fe/H]$, which are taken from the Harris catalog (1996AJ....112.1487H 1996AJ....112.1487H, 2010 edition), all other quantities are derived from the base data provided by the globular cluster database 555people.smp.uq.edu.au/HolgerBaumgardt/globular/666people.smp.uq.edu.au/HolgerBaumgardt/globular/parameter.html. The mass-loss fraction is calculated following the prescription of 2020A&A...637L...2P.
@@ 4.2.1 Possible group-sensitive parameters
$\log_{\rm 10}(M/M_{\odot})$ and $v_{\mathrm{esc}}$: from the analysis of Fig. 6, we observe a positive correlation between $\log_{\rm 10}(M/M_{\odot})$ and $v_{\mathrm{esc}}$. Clusters in Group G1, which have lower mass, also exhibit lower escape velocity, whereas clusters in Group G2 and G3 generally occupy higher values in both parameters. This behavior is consistent with theoretical expectations, as the total mass determines the depth of a cluster’s gravitational potential well and shallower gravitational potential wells make clusters more susceptible to external perturbations and easy to escape (2003MNRAS.340..227B). The two-sample K-S tests reveal significant differences between G1 and G2 ($q=0.02$ for $\log M$ and $q=0.04$ for $v_{\mathrm{esc}}$) as well as between G1 and G3 ($q=0.01$ for $\log M$ and $v_{\mathrm{esc}}$). In contrast, no significant differences are found between G2 and G3 ($q>0.1$). These results show that G1 clusters are clearly distinct from the other two groups. The absence of significant differences between G2 and G3 suggests that G2 clusters may occupy an intermediate state rather than forming a fully separate class. In this sense, G2 may represent a transitional population.
@@
$c=\log(r_{\rm t}/r_{\rm c})$: this parameter reflects the central compactness of the cluster. Examining the distribution of $c$ across the groups, we find a pattern similar to that of $\log_{\rm 10}(M/M_{\odot})$ and $v_{\mathrm{esc}}$: clusters in G1 predominantly occupy lower values, whereas those in G2 and G3 are concentrated at higher values. This trend is consistent with theoretical expectations. 1990ApJ...351..121C found that clusters with lower central concentration and flatter mass functions are more easily disrupted due to cluster expansion driven by stellar evolution and tidal mass loss. Therefore, more compact clusters are likely less susceptible to developing tidal structures. Moreover, Fig. 6 reveals clear positive correlations between $\log_{\rm 10}(M/M_{\odot})$ and $c$, as well as between $v_{\mathrm{esc}}$ and $c$, indicating that more massive clusters generally have higher central concentrations and stronger gravitational binding, making them less prone to forming tidal tails. Two-sample K-S tests further indicate significant differences between G1 and G2/G3 ($q=0.01$), while G2 and G3 show similar distributions ($q=0.27$), supporting the interpretation of G2 as a transitional population.
@@
$r_{\rm h,m}/r_{\rm t}^{\rm dyn}$: defined as the ratio between the half-mass radius and the dynamical tidal radius, this factor measures how tidally filled a cluster is. From Fig. 6, the distribution of this parameter in G1 differs from that in G2/G3: G1 spans the entire parameter space, whereas G2/G3 are confined to lower values. The K-S tests further confirm a significant difference between G1 and G3 ($q=0.04$), while no significant difference is found between G1 and G2 ($q=0.09$) or between G2 and G3 ($q=0.18$). In addition, the correlation plots show that this parameter is noticeably anti-correlated with $\log_{\rm 10}(M/M_{\odot})$, $V_{\mathrm{esc}}$, and $c$. This can be explained theoretically, since more massive or more compact clusters are more resistant to tidal filling and subsequent disruption (1990ApJ...351..121C).
@@
$M_{\rm dis}/M_{\rm init}$: this ratio quantifies the fractional mass loss from the cluster’s initial mass to its current mass. From Fig. 6, we also observe a clear distributional difference between G1 and G3: Clusters in G1 tend to lose a larger fraction of their member stars, making them more prone to developing tidal structures, whereas those in G3 lose fewer stars and are therefore less likely to form such features. G2 appears to represent a transitional population between the two. This parameter also exhibits weak correlations with several physical quantities: clusters with higher $\log_{\rm 10}(M/M_{\odot})$, larger $v_{\mathrm{esc}}$, or higher $c$ tend to show lower $M_{\rm dis}/M_{\rm init}$, whereas more tidally filled systems generally display higher values of $M_{\rm dis}/M_{\rm init}$. In addition, its positive correlation with $\log(\mathrm{age}/T_{\rm rh})$, suggests that clusters that have lost more stars also experience more dynamical relaxation. However, the K-S tests do not reveal any significant differences among the three groups ($q>0.1$). We speculate that this may be influenced by the G2 clusters, as these systems also exhibit rich extra-tidal features and may have experienced a comparable degree of mass loss. In addition, the coupling between this parameter and the clusters’ orbital dynamical properties may introduce further scatter, thereby weakening the statistical separation. Therefore, although this parameter carries meaningful physical correlations, we temporarily classify it as group-insensitive.
@@
$\rm[Fe/H]$: this quantity is widely used as a tracer of the formation environment and epoch, and may also indicate whether a cluster originated in an external satellite galaxy. Metal-poor clusters are often linked to accreted populations and may have undergone more complex interaction histories (2019A&A...630L...4M; 2020MNRAS.493..847F), potentially shaping their outer structures. In our analysis, however, we do not find significant differences in its distribution across different groups. 2025AJ....170..157K recently reported a weak correlation between metallicity and tidal morphology, but their sample is restricted to relatively nearby and massive clusters, which could bias the inferred relation.
@@
$\log_{\rm 10}(\mathrm{age}/T_{\rm rh})$: the logarithmic ratio between the cluster’s age and its half-mass relaxation time indicates the degree of dynamical relaxation. From the parameter distributions in Fig. 6, there is no obvious difference among the groups. The pairwise correlation plots reveal the previously mentioned correlation between $\log(\mathrm{age}/T_{\rm rh})$ and $M_{\rm dis}/M_{\rm init}$. In the study by 2019MNRAS.489.4367P, it was proposed that globular clusters with relatively short relaxation times often have already undergone substantial mass loss. Such depletion, typically caused by tidal stripping or evaporation, weakens the system’s gravitational binding and accelerates its internal dynamical evolution, thereby increasing the likelihood of further stellar escape. This interpretation is consistent with both theoretical and numerical studies: once mass loss begins, the relaxation time shortens, which enhances the efficiency of two-body interactions and accelerates core evolution. For example, 2010MNRAS.409..305L pointed out that external tidal perturbations and the gradual depletion of loosely bound stars can drive a positive feedback loop between mass loss and structural evolution, while 2008MNRAS.389L..28G showed that clusters evolving within the realistic Galactic tidal field exhibit a weaker scaling between dissolution time and relaxation time than in isolation ($t_{\rm dis}\propto t_{\rm rh}^{3/4}$), highlighting the accelerating role of tidal stripping. Thus, a short relaxation time may not only be interpreted as the dynamical consequence of the preferential loss of outer stars, but also as a potential precursor to the emergence of extended tidal features. In our analysis, the parameters $M_{\rm dis}/M_{\rm init}$ and $\log(\mathrm{age}/T_{\rm rh})$ show a certain degree of linear correlation, which may reflect the influence of such evolutionary processes. However, no significant clustering or separation is observed either in this parameter space or in those involving $\log(\mathrm{age}/T_{\rm rh})$. This suggests that while relaxation time and its related ratios play an important role in the dynamical evolution of globular clusters, they alone are insufficient to reliably characterize the presence or extent of tidal structures, particularly given the diverse environments and evolutionary stages represented in our sample.
@@
Table 3: Two-sample Kolmogorov-Smirnov (K-S) test results for internal and orbital parameters among different cluster groups. [^7] Col.1: the parameter name; Col.2-4: results of the pairwise K-S tests between the three groups (G1 vs G2, G1 vs G3, G2 vs G3), reported as the FDR-adjusted $p$-values ($q$).

### v2

@@ 4.1 Detection of tidal structures
Based on the matched filtering technique, we constructed S/N maps for all clusters in our sample (see Fig. 2). The maps are presented in a projected coordinate system centered on each cluster, with axes defined as $\Delta\mathrm{RA}\times\cos(\mathrm{Dec})$ and $\Delta\mathrm{Dec}$, ensuring consistent spatial analysis across different celestial positions. The cluster centers were adopted from the Galactic globular cluster database 333people.smp.uq.edu.au/HolgerBaumgardt/globular/, which is extensively used in this study. This database, developed by Baumgardt and collaborators, provides a comprehensive set of physical and dynamical parameters for Galactic GCs. These parameters were derived by combining a compilation of observational data, including ground-based radial velocities, Gaia DR3 proper motions, and HST-based stellar mass functions, with literature data, and by performing N-body modeling and orbital integrations (see, e.g., Baumgardt 2017; Baumgardt and Hilker 2018; Vasiliev and Baumgardt 2021; Baumgardt and Vasiliev 2021; Baumgardt et al. 2023, etc.). Further details can be found in the database.
@@
We adopted a uniform spatial resolution of $0.02^{\circ}$ ($1\arcmin.2$) for all clusters except for Whiting 1. This cluster is a relatively small system, with a King-model tidal radius of only $0\arcmin.89$ according to the Harris catalog (Harris 1996, 2010 edition), and it lies at a heliocentric distance of $30.59$ kpc (Baumgardt and Vasiliev 2021). As a result, its projected angular extent on the sky is limited. To better resolve its structural features, we adopted a finer spatial resolution of $0.01^{\circ}$ ($0\arcmin.6$) for this target. In addition, we applied Gaussian smoothing to the matched-filter S/N maps to enhance the visibility of extended structures while suppressing small-scale noise. Because the clusters in our sample differ substantially in distance, apparent angular size, and the spatial distribution of their member stars, a single smoothing scale is not appropriate. We therefore adopted $\sigma$ = 1 pixel as the default smoothing kernel. The kernel size was then adjusted through systematic visual inspection of the resulting S/N maps to identify the smallest kernel that effectively reduces noise without oversmoothing genuine features. Such inspections help effectively reduce the risk of smoothing out real features or missing potential tidal structures from insufficient smoothing. In practice, the smoothing kernels adopted for our sample range from $\sigma$=1 to 5 pixels, depending on the cluster’s properties.
@@
To investigate the properties of extratidal features among our sample clusters, we analyzed the matched-filtered S/N maps and adopted a classification scheme similar to that of Piatti and Carballo-Bello (2020). Specifically, we overlaid two reference tidal radii on each map: the King tidal radius (${r_{\rm t}}^{\rm King}$) taken from the Harris catalog (Harris 1996, 2010 edition), which is derived from surface density fitting under the King model (King 1962), and the dynamical tidal radius (${r_{\rm t}}^{\rm dyn}$) from the globular cluster database 444people.smp.uq.edu.au/HolgerBaumgardt/globular/parameter.html, which represents the present-day limiting radius $r_{\rm L}(F)$ calculated following the expression of Webb et al. (2013):
@@
where $R_{\rm GC}$ is the current Galactocentric distance of the cluster, and $R_{\rm Apo}$ and $R_{\rm Peri}$ denote the apocenter (maximum) and pericenter (minimum) distances along its orbit, respectively. The coefficients are $a=0.17\pm 0.03$ and $b=4.1\pm 0.2$ (Webb et al. 2013). Compared to the King tidal radius, the dynamical tidal radius provides a physically motivated boundary that accounts for orbital dynamics and gravitational interactions with the Galactic potential.
All clusters in our sample were classified into three groups according to their structural properties with respect to $r_{\rm t}^{\rm King}$ and ${r_{\rm t}}^{\rm dyn}$: G1 clusters exhibit coherent, tail-like extensions aligned along a preferred direction, with detectable overdensities (typically reaching $\mathrm{S/N}\geq 1\sigma$) that extend beyond the dynamical tidal radius ($r_{\rm s}>r_{\rm t}^{\rm dyn}$); G2 clusters exhibit overdensities with $\mathrm{S/N}\geq 3\sigma$ extending beyond the King tidal radius ($r_{\rm s}>r_{\rm t}^{\rm King}$), while being overall confined within the dynamical tidal radius ($r_{\rm s}<r_{\rm t}^{\rm dyn}$). These extensions typically appear as diffuse envelopes, irregular overdensities, or clumpy halo-like features, and may in some cases extend beyond $r_{\rm t}^{\rm dyn}$. In addition, clusters that display tail-like structures beyond $r_{\rm t}^{\rm King}$ but still remain fully contained within $r_{\rm t}^{\rm dyn}$ are also included in this group; G3 clusters show no overdensities with $\mathrm{S/N}\geq 3\sigma$ beyond the King tidal radius ($r_{\rm s}<r_{\rm t}^{\rm King}$), and whose outer stellar density profiles are consistent with the truncation expected from a King model.
@@ 4.2 Dependence of tidal features on intrinsic properties
To investigate whether any internal parameters can effectively predict the presence or morphology of tidal structures in GCs, we selected a representative set of physical quantities that are commonly associated with the structural and dynamical evolution of clusters. These include the present-day cluster mass ($\log M$), central escape velocity ($v_{\rm esc}$), central concentration ($c$), ratio of half-mass radius to dynamical tidal radius ($r_{\rm h,m}/r_{\rm t}^{\rm dyn}$), mass-loss fraction ($M_{\rm dis}/M_{\rm init}$), metallicity ($\rm[Fe/H]$), and relaxation-age ratio ($\log({\rm age}/{T}_{\rm rh})$). Except for $c$ and $\rm[Fe/H]$, which are taken from the Harris catalog (Harris 1996, 2010 edition), all other quantities are derived from the base data provided by the globular cluster database 555people.smp.uq.edu.au/HolgerBaumgardt/globular/666people.smp.uq.edu.au/HolgerBaumgardt/globular/parameter.html. The mass-loss fraction is calculated following the prescription of Piatti and Carballo-Bello (2020).
@@ 4.2.1 Possible group-sensitive parameters
From analysis of Fig. 3, we observed a positive correlation between $\log_{\rm 10}(M/M_{\odot})$ and $v_{\mathrm{esc}}$. Clusters in Group G1, which have lower mass, also exhibit lower escape velocity, whereas clusters in Group G2 and G3 generally occupy higher values in both parameters. This behavior is consistent with theoretical expectations, as the total mass determines the depth of a cluster’s gravitational potential well and shallower gravitational potential wells make clusters more susceptible to external perturbations and easy to escape (Baumgardt and Makino 2003). The two-sample K-S tests reveal significant differences between G1 and G2 ($q=0.02$ for $\log M$ and $q=0.04$ for $v_{\mathrm{esc}}$) as well as between G1 and G3 ($q=0.01$ for $\log M$ and $v_{\mathrm{esc}}$). In contrast, no significant differences are found between G2 and G3 ($q>0.1$). These results show that G1 clusters are clearly distinct from the other two groups. The absence of significant differences between G2 and G3 suggests that G2 clusters may occupy an intermediate state rather than forming a fully separate class. In this sense, G2 may represent a transitional population.
@@
The parameter $c=\log(r_{\rm t}/r_{\rm c})$ reflects the central compactness of the cluster. Examining the distribution of $c$ across the groups, we find a pattern similar to that of $\log_{\rm 10}(M/M_{\odot})$ and $v_{\mathrm{esc}}$: clusters in G1 predominantly occupy lower values, whereas those in G2 and G3 are concentrated at higher values. This trend is consistent with theoretical expectations. Chernoff and Weinberg (1990) found that clusters with lower central concentration and flatter mass functions are more easily disrupted due to cluster expansion driven by stellar evolution and tidal mass loss. Therefore, more compact clusters are likely less susceptible to developing tidal structures. Moreover, Fig. 3 reveals clear positive correlations between $\log_{\rm 10}(M/M_{\odot})$ and $c$, as well as between $v_{\mathrm{esc}}$ and $c$, indicating that more massive clusters generally have higher central concentrations and stronger gravitational binding, making them less prone to forming tidal tails. Two-sample K-S tests further indicate significant differences between G1 and G2/G3 ($q=0.01$), while G2 and G3 show similar distributions ($q=0.27$), supporting the interpretation of G2 as a transitional population.
@@
Defined as the ratio between the half-mass radius and the dynamical tidal radius, $r_{\rm h,m}/r_{\rm t}^{\rm dyn}$ measures how tidally filled a cluster is. From Fig. 3, the distribution of this parameter in G1 differs from that in G2/G3: G1 spans the entire parameter space, whereas G2/G3 are confined to lower values. The K-S tests further confirm a significant difference between G1 and G3 ($q=0.04$), while no significant difference is found between G1 and G2 ($q=0.09$) or between G2 and G3 ($q=0.18$). In addition, the correlation plots show that this parameter is noticeably anticorrelated with $\log_{\rm 10}(M/M_{\odot})$, $v_{\mathrm{esc}}$, and $c$. This can be explained theoretically, since more massive or more compact clusters are more resistant to tidal filling and subsequent disruption (Chernoff and Weinberg 1990).
@@ 4.2.2 Group-insensitive parameters
The ratio $M_{\rm dis}/M_{\rm init}$ quantifies the fractional mass loss from the cluster’s initial mass to its current mass. From Fig. 3, we also observe a clear distributional difference between G1 and G3: Clusters in G1 tend to lose a larger fraction of their member stars, making them more prone to developing tidal structures, whereas those in G3 lose fewer stars and are therefore less likely to form such features. G2 appears to represent a transitional population between the two. This parameter also exhibits weak correlations with several physical quantities: clusters with higher $\log_{\rm 10}(M/M_{\odot})$, larger $v_{\mathrm{esc}}$, or higher $c$ tend to show lower $M_{\rm dis}/M_{\rm init}$, whereas more tidally filled systems generally display higher values of $M_{\rm dis}/M_{\rm init}$. In addition, its positive correlation with $\log(\mathrm{age}/T_{\rm rh})$, suggests that clusters that have lost more stars also experience more dynamical relaxation. However, the K-S tests do not reveal any significant differences among the three groups ($q>0.1$). We speculate that this may be influenced by the G2 clusters, as these systems also exhibit rich extratidal features and may have experienced a comparable degree of mass loss. In addition, the coupling between this parameter and the clusters’ orbital dynamical properties may introduce further scatter, thereby weakening the statistical separation. Therefore, although this parameter carries meaningful physical correlations, we temporarily classify it as group-insensitive.
@@
The quantity $\rm[Fe/H]$ is widely used as a tracer of the formation environment and epoch, and may also indicate whether a cluster originated in an external satellite galaxy. Metal-poor clusters are often linked to accreted populations and may have undergone more complex interaction histories (Massari et al. 2019; Forbes 2020), potentially shaping their outer structures. In our analysis, however, we do not find significant differences in its distribution across different groups. Kuzma et al. (2025) recently reported a weak correlation between metallicity and tidal morphology, but their sample is restricted to relatively nearby and massive clusters, which could bias the inferred relation.
@@
The logarithmic ratio $\log_{\rm 10}(\mathrm{age}/T_{\rm rh})$ between the cluster’s age and its half-mass relaxation time indicates the degree of dynamical relaxation. From the parameter distributions in Fig. 3, there is no obvious difference among the groups. The pairwise correlation plots reveal the previously mentioned correlation between $\log(\mathrm{age}/T_{\rm rh})$ and $M_{\rm dis}/M_{\rm init}$. In the study by Piatti et al. (2019), it was proposed that GCs with relatively short relaxation times often have already undergone substantial mass loss. Such depletion, typically caused by tidal stripping or evaporation, weakens the system’s gravitational binding and accelerates its internal dynamical evolution, thereby increasing the likelihood of further stellar escape. This interpretation is consistent with both theoretical and numerical studies: once mass loss begins, the relaxation time shortens, which enhances the efficiency of two-body interactions and accelerates core evolution. For example, Lamers et al. (2010) pointed out that external tidal perturbations and the gradual depletion of loosely bound stars can drive a positive feedback loop between mass loss and structural evolution, while Gieles and Baumgardt (2008) showed that clusters evolving within the realistic Galactic tidal field exhibit a weaker scaling between dissolution time and relaxation time than in isolation ($t_{\rm dis}\propto t_{\rm rh}^{3/4}$), highlighting the accelerating role of tidal stripping. Thus, a short relaxation time may not only be interpreted as the dynamical consequence of the preferential loss of outer stars, but also as a potential precursor to the emergence of extended tidal features. In our analysis, the parameters $M_{\rm dis}/M_{\rm init}$ and $\log(\mathrm{age}/T_{\rm rh})$ show a certain degree of linear correlation, which may reflect the influence of such evolutionary processes. However, no significant clustering or separation is observed either in this parameter space or in those involving $\log(\mathrm{age}/T_{\rm rh})$. This suggests that while relaxation time and its related ratios play an important role in the dynamical evolution of GCs, they alone are insufficient to reliably characterize the presence or extent of tidal structures, particularly given the diverse environments and evolutionary stages represented in our sample.
@@
Table 3: Two-sample K-S test results for internal and orbital parameters among different cluster groups. [^7] Col.1: parameter name; Col.2-4: results of the pairwise K-S tests between the three groups (G1 vs. G2, G1 vs. G3, G2 vs. G3), reported as the FDR-adjusted $p$-values ($q$).